OpenStudy (romantic_ch3micals):

Hi, I need help doing these two. Can someone explain. Medal How many real number solutions does the equations have? 0 = 5x^2 + 2x - 12 0 = 2x^2 - 20x + 50

OpenStudy (campbell_st):

do you know about the discriminant \[\Delta = b^2 - 4ac\]

OpenStudy (dtan5457):

Yup ^, all you need is the discriminant. If it equals 0, it's two equal real solutions. <0, 2 imaginary solutions >0, rational or irrational, equal or unequal depending if it's a perfect square

OpenStudy (romantic_ch3micals):

So they both have two equal solutions because they equal 0? @dtan5457

OpenStudy (dtan5457):

My bad, if it's equal to 0, there is one real solution.

OpenStudy (romantic_ch3micals):

Oh, well, I get that but is there something else I should know? I really want to get this.

OpenStudy (campbell_st):

= 0 its 1 repeated real solution

OpenStudy (romantic_ch3micals):

Ok, so they both have one solution because it equals to 0. Am I correct?

OpenStudy (dtan5457):

Are you referring to your questions?

OpenStudy (romantic_ch3micals):

Yeah.

OpenStudy (dtan5457):

You need to use the discriminant of b^2-4ac to figure that out. Just because the quadratic=0, doesn't mean the discriminant is 0.

OpenStudy (dtan5457):

I'll do the first one. 0 = 5x^2 + 2x - 12 B^2-4ac value of b=2 value of a=5 value of c=-12 (2)^2-4(5)(-12)=244 244 is not a perfect square. therefore it has 2 unequal, irrational real solutions. try the second one.

OpenStudy (romantic_ch3micals):

Ok. eh, Ive work it out on paper and I got three...is that right? @dtan5457